The tens digit of a two-digit number is half the units digit. If their difference is 4, what is the number?
Answer by stanbon(48575)

The tens digit of a two-digit number is half the units digit. If their difference is 4, what is the number?
----------------------
Let the number be 10t+u.
-----------
Equations:
t = (1/2)u
u -t = 4
--------------------
Substitute to solve for "u":
u - (1/2)u = 4
(1/2)u = 4
u = 8
----
Substitute to solve for "t":
t = (1/2)u = 4
--------------------
Ans: The number is 48
==========================
Cheers,
Stan H.